# Imperceptible Protection against Style Imitation from Diffusion Models

### v2

@@ Introduction
The groundbreaking advancements in large-scale diffusion models have transformed media creation workflows (Rombach et al. 2022; Balaji et al. 2022; Li et al. 2022; Podell et al. 2023; Saharia et al. 2022). These can be further enhanced in usability when integrated with external modules that accept multi-modal inputs (Zhang, Rao, and Agrawala 2023; Li et al. 2023; Mou et al. 2023; Huang et al. 2023; Kim et al. 2023). Such innovations have also been pivotal in the realm of art creation (Ko et al. 2023; Sohn et al. 2023; Ahn et al. 2023). Nevertheless, generative AI, while undoubtedly beneficial, brings concerns about its potential misuse. When someone exploits these to replicate artworks without permission, it introduces significant risks of copyright infringement. This style imitation becomes a serious threat to artists (Shan et al. 2023; Xiang 2022).
To counteract style imitation, previous studies have introduced adversarial perturbation (Goodfellow, Shlens, and Szegedy 2014; Madry et al. 2017) to artwork, transforming it into an adversarial example that can resist few-shot generation or personalization methods (Van Le et al. 2023; Liang et al. 2023; Liang and Wu 2023; Ye et al. 2023; Zheng et al. 2023; Salman et al. 2023; Zhao et al. 2023). Specifically, building on the Stable Diffusion (SD) model (Rombach et al. 2022), they iteratively optimize the protected (or perturbed) image to fool the diffusion models, guided by the gradients from the image encoder or denoising UNet.
@@
While existing studies are effective in preventing style imitation, they do not prioritize the protected image’s quality. They leave discernible traces (or artifacts) on the protected images due to the inherent nature of adversarial perturbations. Moreover, we observed that unlike adversarial attacks on classifiers (Luo et al. 2022; Dai et al. 2023), style protection requires more intense and globally dispersed perturbations. Consequently, despite the commendable protection performance, prior works run the risk of severely degrading the original artwork’s fidelity, making it less practical for real-world applications. While moderating the strength of protection could mitigate this, it introduces a trade-off, often compromising protection performance and making it challenging to achieve satisfactory results.
@@
To address this, we propose Impasto (Figure 1). We design upon the principle of perception-aware protection, which focuses on perturbing regions less discernible to humans. While many studies in the adversarial attacks regime limit perturbations to small areas to maximize imperceptibility (Croce and Hein 2019; Dai et al. 2023), they are ineffective for style protection, as personalization methods can leverage references from non-perturbed textures. We instead adopt a soft restriction strategy, which protect the entire image but with modulated intensities, relaxing the harsh condition of sparse constraints. To implement this, it is crucial to identify which areas are perceptually more noticeable when perturbations are introduced. For this purpose, we analyze various perceptual maps that are suitable for our task. Then, we propose a method that combines such perceptual maps and refines them in an image-specific manner. Such an instance-wise refinement sets itself apart from previous protection methods as it offers a way to strike the optimal balance between imperceptibility and protection performance.
@@
For better imperceptibility, we propose a difficulty-aware protection that automatically adjusts the protection strength based on how easily the area can be protected. To achieve this, we compute a difficulty map, which predicts the protection difficulty of a given artwork. The difficulty map is then combined with the perceptual map to determine the final protection intensity. Lastly, we introduce a perceptual constraint bank, which uses various features to apply perceptual constraints. It may seem unsurprising that a constraint bank can improve imperceptibility intuitively. However, our study is the first to apply constraints across multiple spaces, improving fidelity without compromising robustness.
@@
Figure 1: Model overview. (a) Previous methods protect image through an iterative process that updates perturbations $\delta$ with PGD. (b) On this basis, Impasto constructs a perceptual map $\mathcal{M}$ and a difficulty map $\mathcal{M_{D}}$ to adaptively refine the perturbation for imperceptibility. Additionally, Impasto employs a constraint bank in the PGD process to achieve better imperceptibility. Red-bordered boxes indicate trainable parameters. IWR denotes instance-wise refinement.
@@
To the best of our knowledge, Impasto is the pioneering approach that prioritizes the protected image’s quality in the style protection task. It achieves robust protection against style imitation and maintains visual fidelity in protected images thus it can significantly improve the trade-off balance between protection efficacy and image quality. Moreover, Impasto is versatile, enabling its attachment to any existing protection frameworks (Liang et al. 2023; Salman et al. 2023; Van Le et al. 2023; Liang and Wu 2023). Impasto also maintains resilience and generalizes well against a range of countermeasures and personalization techniques, performing on par with the baseline methods.
Background
@@
Diffusion Models. These have gain prominence for their capacity to produce high-quality images (Ho, Jain, and Abbeel 2020; Dhariwal and Nichol 2021; Saharia et al. 2022; Nichol et al. 2021; Balaji et al. 2022). In AI-assisted art production, Stable Diffusion (SD) (Rombach et al. 2022) is widely used because of its efficiency. In SD, an input image $\mathbf{x}$ is projected into a latent code via an image encoder $\mathcal{E}$; $\mathbf{z}=\mathcal{E}(\mathbf{x})$. A decoder $\mathcal{D}$ reverts the latent code to the image domain; $\mathbf{x}^{\prime}=\mathcal{D}(\mathbf{z}^{\prime})$. The diffusion model derives a modified code $\mathbf{z}^{\prime}$ using external factors $y$, such as text prompt or other modalities (Zhang, Rao, and Agrawala 2023; Kim et al. 2023). The training objective for SD at timestep $t$ is defined as:
$\mathcal{L_{SD}}=\mathbb{E}_{\mathbf{z}\sim E(\mathbf{x}),y,\epsilon\sim N(0,1),t}[||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,c(y))||^{2}_{2}].$ (1)
@@
(a) No protect (b) Partial protect (c) Full protect
Figure 2: Comparison of restriction types. We compare partial and full protection by fine-tuning diffusion models using protected (or original) images via DreamBooth (Ruiz et al. 2023). For partial protection, perturbations are applied to the center (or facial) region. The results indicate that partial protection is insufficient for style protection, as personalization methods can utilize unprotected areas to learn the artwork’s style. Reference images are shown in Suppl.
@@
Figure 3: Examples of perceptual maps. In this study, we use luminance adaptation (LA), contrast masking (CM), contrast sensitivity function (CSF), standard deviation (Stdev), and entropy. (Left) Perceptual Map visualization. Darker region corresponds to increased protection intensity. Impasto constructs perceptual map $\mathcal{M}$ by spatially averaging these estimations (Avgerage) or an learnable manner (IWR). (Right) Perceptual map comparison. We evaluate the protection performance of each perceptual map. Image quality is assessed through DISTS (Ding et al. 2020), and protection performance is evaluated via FID (Heusel et al. 2017). The baseline is trained with Eq. 5 and other models are trained via Eq. 6 with corresponding JNDs.
Protection Against Style Imitation. Previous methods introduce adversarial perturbations $\delta$ to image $\mathbf{x}$, making protected image $\hat{\mathbf{x}}=\mathbf{x}+\delta$ through projected gradient descent (PGD) (Madry et al. 2017). 1) Encoder-based methods (Shan et al. 2023; Salman et al. 2023; Ye et al. 2023; Liang and Wu 2023) (e.g. PhotoGuard) update $\delta$ under the guidance of the encoder $\mathcal{E}$. They aim to maximize the distance between the feature of the original and protected image. In practice, many methods minimize the distance between the protected image and target image $\mathbf{y}$ instead:
$\delta=\operatorname*{arg\,min}_{||\delta||_{\infty}\leq\eta}\,\mathcal{L_{E}}(\mathbf{x}+\delta,\;\mathbf{y}),\;\;\mathcal{L_{E}}=||\mathcal{E}(\mathbf{x}+\delta)-\mathcal{E}(\mathbf{y})||^{2}_{2}.$ (2)
While $L_{\infty}$ norm ($||\delta||_{\infty}\leq\eta$;  $\eta$ is a budget) is widely used, GLAZE (Shan et al. 2023) adopts LPIPS (Zhang et al. 2018) and DUAW (Ye et al. 2023) employs SSIM (Wang et al. 2004). 2) Diffusion-based methods (Van Le et al. 2023; Liang et al. 2023; Salman et al. 2023; Zhao et al. 2023; Liang and Wu 2023) (e.g. AdvDM) update $\delta$ under the guidance of UNet, $\epsilon_{\theta}$, maximizing diffusion loss $\mathcal{L_{SD}}$:
$\delta=\operatorname*{arg\,max}_{||\delta||_{\infty}\leq\eta}\,\mathcal{L_{SD}}(\mathcal{E}(\mathbf{x}+\delta)).$ (3)
Upon this, Anti-DreamBooth (Van Le et al. 2023) integrates DreamBooth training and Mist (Liang and Wu 2023) merges Eq. 2 and 3, enhancing performance and robustness in various scenarios. Since Impasto is versatile and can be effortlessly integrated into any existing protection methods, we generalize the style protection as below formulation.
$\delta=\operatorname*{arg\,max}_{||\delta||_{\infty}\leq\eta}\,\mathcal{L_{SP}}(\mathbf{x}+\delta,\mathbf{y}),$ (4)
Method
Perception-Aware Protection
Naive sparse restriction. In adversarial attacks, perturbations are often confined to a sparse region to increase imperceptibility (Croce and Hein 2019; Dai et al. 2023). Yet, we observed that they can not sufficiently prevent style imitation. In Figure 2, we compare the protection encompassing the entire image against the one that applies perturbation to facial region only. It is shown that the partial protection cannot protect original artwork against DreamBooth (Ruiz et al. 2023) and we argue that this is because the personalization method can leverage textures from unprotected regions. Overall, sparse restriction may confine the perturbed region too aggressively, thus we relax such assumptions to better align with the style protection task.
@@
Perceptual map. To build a perceptual map $\mathcal{M}$ in a simple yet effective manner, we investigate the just noticeable difference (JND) concept (Wu, Shi, and Lin 2019) inspired by the human visual system. JND represents the minimum intensity of stimulus (perturbation in our context) required to produce a noticeable change in visual perception. The intent behind the JND estimation model is to determine this perceptual threshold. Given that the fundamental premise of JND— to quantify human sensitivity to subtle changes— aligns with our objective of perception-aware protection, we focus on analyzing the effectiveness of JNDs in our formulation, and which JND models yield the good results.
To this end, as shown in Figure 3, we compare five widely used JNDs: 1) luminance adaptation (LM), which considers pixel luminance (Jarsky et al. 2011), 2) contrast masking (CM), which measures scene complexity by luminance contrast (Legge and Foley 1980; Wu et al. 2013a), 3) contrast sensitivity function (CSF), which utilizes frequency signals (Wei and Ngan 2009), 4) standard deviation (Stdev), which measures spatial structure with patch-wise statistics, and 5) entropy, which computes complexity with patch-wise entropy. More details are described in Suppl.
In our observation, most JND models fit surprisingly well with soft restriction (Figure 3, right). Consequently, we employ a JND-based perceptual maps, considering its simplicity and effectiveness. Note that among JNDs we used, LA and CM emerge as superior in the quality-protection trade-off. However, these results are the average scores across a dataset and we note that some artworks exhibit better trade-off with other JNDs. Especially, since artworks have diverse styles, the best combinations can differ; for example, in Figure 3 (left), Stdev and entropy are best for top image while bottom image shows a preference for CM and CSF. Moreover, in practical scenarios, users who protect their artwork are solely concerned with their specific pieces, not the average scores. Thus, it is crucial to ensure that the protection method is effectively applied to every individual artwork.
@@
Instance-wise refinement. Although the above method is more effective in many scenarios than the single JND, applying a uniform averaged map across all artworks may still be suboptimal for some artworks. Moreover, the optimal perceptual map corresponding to an artwork varies depending on both artwork’s structure and the applied protective perturbations. Even for identical artwork, the detectability of perturbations is affected by changing these as human sensitivity varies with distortion type (Dodge and Karam 2017). Considering these, we propose an instance-wise refinement (IWR), customizing the perceptual map $\mathcal{M}$ for each artwork. During optimization steps, the perceptual map is refined through a weighted sum: $\mathcal{M}(\omega)=\textstyle\sum_{k=1}^{K}\text{softmax}(\mathbf{\omega})^{k}*M^{k}$, where $\mathbf{\omega}$ is a set of learnable parameters that adjust the contributions of each JND to $\mathcal{M}$. The refinement parameters $\mathbf{\omega}$ is optimized using the objective function below:
$\mathcal{L_{M}}$ $=||\mathcal{L_{SP}}(\mathbf{x}+\mathcal{M}^{\prime}\odot\delta)-\mathcal{L_{SP}}(\mathbf{x}+\mathcal{M}(\omega)\odot\delta)||^{2}_{2}$
@@
with $\mathcal{M}^{\prime}$ being the initial perceptual map before the refinement steps. In Eq. 7, the former term enforces the consistency of the refined perceptual map $\mathcal{M}$ with the initial map $\mathcal{M}^{\prime}$ by minimizing the discrepancy in protection loss between them. The latter term enhances the perception-aware protection for a given specific image. As shown in Figure 3, instance-wise refinement (IWR) has a pronounced effect as compared to the naive averaging, better capturing the nuances of image-specific perceptual sensitivity.
@@
Figure 4: Protection difficulty. We present the protection efficacy of PhotoGuard (Salman et al. 2023) across 20 artworks (x-axis). Notably, the level of style protection varies widely between artworks, thus the average score (red dashed line) is less meaningful to users who protect their artwork.

### v4

@@ I Introduction
The groundbreaking advancements in large-scale diffusion models have transformed media creation workflows [47, 2, 27, 44, 49]. These can be further enhanced in usability when integrated with external modules that accept multi-modal inputs [69, 28, 41, 19, 22]. Such innovations have also been pivotal in the realm of art creation [23, 56, 1]. Nevertheless, generative AI, while undoubtedly beneficial, brings concerns about its potential misuse. When someone exploits these to replicate artworks without permission, it introduces significant risks of copyright infringement. This style imitation becomes a serious threat to artists [53, 66]. In addition, a recent survey by the Society of Authors found that 37% of UK illustrators saw their primary‑sale income fall by 12–18% after their styles were cloned by generative models, and 26% had already lost work because of generative AI [55]. By preventing unauthorized style cloning, creators in the US and EU are more likely to retain clients who might otherwise turn to inexpensive AI-generated imitations. Meanwhile, popular IPs in Japan and Korea (e.g., manga and webtoon styles) may remain exclusive to their original artists. This also includes the protection of comic book sales, a market projected to reach approximately \$17.5 billion by 2030 [14].
@@
To counteract style imitation, previous studies have introduced adversarial perturbation [13, 33] to artwork, transforming it into an adversarial example that can resist few-shot generation or personalization methods [59, 29, 30, 68, 73, 50, 72]. Specifically, building on the Stable Diffusion (SD) model [47], they iteratively optimize the protected (or perturbed) image to fool the diffusion models, guided by the gradients from the image encoder or denoising UNet. While existing studies are effective in preventing style imitation, they do not prioritize the protected image’s quality. They leave discernible traces (or artifacts) on the protected images due to the inherent nature of adversarial perturbations. Moreover, we observed that unlike adversarial attacks on classifiers [32, 7], style protection requires more intense and globally dispersed perturbations. Consequently, despite the commendable protection performance, prior works run the risk of severely degrading the original artwork’s fidelity, making it less practical for real-world applications. While moderating the strength of protection could mitigate this, it introduces a trade-off, often compromising protection performance and making it challenging to achieve satisfactory results.
@@
To address this, we propose Impasto (Fig. 1). We design upon the principle of perception-aware protection, which focuses on perturbing regions less discernible to humans. While many studies in the adversarial attacks regime limit perturbations to small areas to maximize imperceptibility [6, 7], they are ineffective for style protection, as personalization methods can leverage references from non-perturbed textures. We instead adopt a soft restriction strategy, which protect the entire image but with modulated intensities, relaxing the harsh condition of sparse constraints. To implement this, it is crucial to identify which areas are perceptually more noticeable when perturbations are introduced. For this purpose, we analyze various perceptual maps that are suitable for our task.
With a given perceptual map, the most suitable map may vary for each input image, hence a one-size-fits-all approach inherently limited in terms of imperceptibility. We instead propose a method that combines multiple perceptual maps and refines them in an image-specific manner. The instance-wise refinement differentiates itself from previous protection methods by striking an optimal balance between imperceptibility and protection performance. Our approach, which gradually modulates image perturbations through a soft restriction mechanism, has been explored in adversarial attack research [40, 33]. However, we are the first to analyze restriction mechanisms in the context of image protection. Furthermore, we introduce an adaptive constraint that optimizes imperceptibility on a per-image basis, a novel contribution to the field.
For better imperceptibility, we propose a difficulty-aware protection, which automatically adjusts protection strength based on how easily each region can be protected. To achieve this, we compute a difficulty map that predicts the protection difficulty of a given artwork. Ideally, evaluating an image’s protection difficulty would require applying protection, performing a personalization method, and measuring how well the style image is mimicked. However, this process is computationally expensive, as it involves both perturbation optimization and personalization. To make this more practical, an approximation is necessary. Our findings suggest that passing a protected image through a diffusion process [17, 35] provides a reasonable approximation. This is because diffusion process tends to amplify injected perturbations, producing patterns similar to those observed in personalization methods. By applying the diffusion process to both the original and protected images and measuring their perceptual distance, we estimate the difficulty map. The difficulty map is then combined with the perceptual map to determine the final protection intensity.
@@
Lastly, we introduce a perceptual constraint bank, which uses various features to apply perceptual constraints. It may seem unsurprising that a constraint bank can improve imperceptibility intuitively. However, our study is the first to apply constraints across multiple spaces, improving fidelity without compromising robustness. In addition, the perceptual constraint bank can be seen as an ensemble of multiple latent space constraints. This approach is more effective than applying constraints solely in the pixel space.
@@
Figure 1: Model overview. (a) Current protection methods update perturbation $\delta$ with an iterative process. (b) On this basis, Impasto constructs a perceptual map $\mathcal{M}$ and a difficulty map $\mathcal{M_{D}}$ to adaptively refine the perturbation for imperceptibility. Impasto also employs a constraint bank in the process to achieve better imperceptibility. Red-bordered boxes indicate trainable parameters. IWR denotes instance-wise refinement.
@@
To the best of our knowledge, Impasto is the pioneering approach that prioritizes the protected image’s quality in the style protection task. It achieves robust protection against style imitation and maintains visual fidelity in protected images thus it can significantly improve the trade-off balance between protection efficacy and image quality. Moreover, Impasto is versatile, enabling its attachment to any existing protection frameworks [29, 50, 59, 30]. Impasto also maintains resilience and generalizes well against a range of countermeasures and personalization techniques, performing on par with the baseline methods.
II Background
@@
Diffusion Models. These have gain prominence for their capacity to produce high-quality images [17, 9, 49, 42, 2]. In AI-assisted art production, Stable Diffusion (SD) [47] is widely used because of its efficiency. In SD, an input image $\mathbf{x}$ is projected into a latent code via an image encoder $\mathcal{E}$; $\mathbf{z}=\mathcal{E}(\mathbf{x})$. A decoder $\mathcal{D}$ reverts the latent code to the image domain; $\mathbf{x}^{\prime}=\mathcal{D}(\mathbf{z}^{\prime})$. The diffusion model derives a modified code $\mathbf{z}^{\prime}$ using external factors $y$, such as text or other modalities [69, 22]. The training objective for SD at timestep $t$ is defined as:
[…]
@@
Protection Against Style Imitation. Previous methods introduce adversarial perturbations $\delta$ to image $\mathbf{x}$, making protected image $\hat{\mathbf{x}}=\mathbf{x}+\delta$ through projected gradient descent (PGD) [33]. 1) Encoder-based methods [53, 50, 68, 30] (e.g. PhotoGuard [50] and GLAZE [53]) update $\delta$ under the guidance of the encoder $\mathcal{E}$. They aim to maximize the distance between the feature of the original and protected image. In practice, many methods minimize the distance between the protected image and target image $\mathbf{y}$ instead:
[…]
While $L_{\infty}$ norm ($\delta||_{\infty}\leq\eta$;  $\eta$ is a budget) is widely used, GLAZE [53] adopts LPIPS [70] and DUAW [68] employs SSIM [61]. 2) Diffusion-based methods [59, 29, 50, 72, 30] (e.g. AdvDM [29]) update $\delta$ under the guidance of UNet, $\epsilon_{\theta}$, maximizing diffusion loss $\mathcal{L_{SD}}$:
[…]
Upon this, Anti-DreamBooth [59] integrates DreamBooth [48] training in the protection optimization while Mist [30] merges Eq. 2 and 3, enhancing performance and robustness in various scenarios. Diff-Protect [67], building on Mist, introduces a score distillation trick to achieve more efficient protection. Since Impasto is versatile and can be effortlessly integrated into any existing protection methods, we generalize the style protection as below formulation.
[…]
Imperceptible adversarial examples. The concept of imperceptibility has been actively investigated in adversarial attacks [31, 15, 54, 51, 39, 6, 52, 71, 60, 32, 7]. Among them, some studies target specific elements; e.g. the low-frequency components [15, 32]. Another approach use advanced constraints [51]; color distance [71] or quality assessment [60] are adopted. Several methods focus on restricting perturbation regions; leveraging $L_{0}$ norm to produce sparse perturbation [39, 6] or limiting perturbations to tiny salient regions [7]. However, we target generative models and uniquely integrates perceptual guidance with instance-wise adaptation, which has not been explored in prior soft restriction approaches [40, 53]. In addition, attacking discriminative models is fundamentally distinct from that of targeting generative models. Hence, we employ a specialized strategy for disrupting style imitation.
Example 1 Refer to caption Refer to caption Refer to caption Refer to caption
Example 2 Refer to caption Refer to caption Refer to caption Refer to caption
(a) Original (b) No protect (c) Partial protect (d) Full protect
Figure 2: Comparison of restriction types. (a) Original image. The red box shows the region with applied partial protection. (b–d) We compare partial and full protection by fine-tuning diffusion models on protected (or original) images with DreamBooth [48]. In the partial protection setting, perturbations are applied only to the central (or facial) area. The results show that partial protection is insufficient for style preservation, as personalization methods can still exploit unprotected regions to capture the artwork’s style.
@@
Figure 3: Examples of perceptual maps. In this study, we use luminance adaptation (LA), contrast masking (CM), contrast sensitivity function (CSF), standard deviation (Stdev), and entropy. (Left) Perceptual Map visualization. Darker region corresponds to increased protection intensity. Impasto constructs perceptual map $\mathcal{M}$ by spatially averaging these estimations (Avgerage) or an learnable manner (IWR). (Right) Perceptual map comparison. We evaluate the protection performance of each perceptual map. Image quality is assessed through DISTS [10], and protection performance is evaluated via FID [16]. The baseline is trained with Eq. 5 and other models are trained via Eq. 6 with corresponding JNDs.
III Method
Impasto aims to enhance the imperceptibility of current image protection methods. Unlike existing approaches that rely on multi-step PGD optimization (Fig. 1a), Impasto integrates multiple imperceptibility-focused modules, allowing it to be versatile and compatible with any protection framework. As shown in Fig. 1b, Impasto consists of three core components: perception-aware protection, difficulty-aware protection, and the constraint bank.
These three modules constitute successive stages of a single pipeline that cooperatively achieve imperceptible yet robust protection: Perception‑Aware Protection (PAP) first decides where to perturb by aligning noise with the human visual system, applying stronger perturbations in regions less noticeable to viewers (§III-A). Difficulty‑Aware Protection (DAP) then calibrates how strongly to perturb each image by predicting its protection difficulty and adapting the global magnitude accordingly (§III-B). Finally, the Perceptual Constraint Bank regularises the optimization trajectory in multiple feature spaces such as texture (with masked LPIPS), color‑frequency (with masked Low‑Pass), and semantics (with CLIP). This provide our method the perturbations selected by PAP + DAP remain visually natural under diverse perceptual lenses (§III-C). Together, this cascade transforms a simple PGD loop into a perception‑ and content‑adaptive procedure.
III-A Perception-Aware Protection
Human observers are markedly more sensitive to noise in smooth, low-texture areas than in highly textured regions. To ensure imperceptible protection, our approach aims to identify regions where noise is least likely to be perceived and to modulate the perturbation strength accordingly on a pixel-wise basis. We start from a straightforward yet insufficient idea of sparsely perturbing only a few pixels and gradually refine it into a soft, map-based restriction that adapts to the visual characteristics of each individual artwork. Figure 2 illustrates the progression of this design.
Naive sparse restriction. In adversarial attacks, perturbations are often limited to sparse regions to enhance imperceptibility [6, 7]. However, we found that this approach does not effectively prevent style imitation. In Fig. 2, we compare full-image protection (Full protect) with protection that applies perturbations only to the facial region (Partial protect). The results show that partial protection fails to safeguard original artwork against DreamBooth [48]. We also perform a quantitative comparison of partial and full protection using FID on a painting dataset. In this test, partial protection yields an FID score of 260.2, while full protection achieves 299.1, demonstrating the superior effectiveness of full protection. These results reveal that simply “hiding” perturbations in a few pixels leaves large unprotected textures that diffusion-based personalization methods can still exploit. We therefore move from an all‑or‑nothing sparse mask to a soft restriction that redistributes perturbation strength continuously across the entire canvas, biased by perceptual visibility.
@@ III-A Perception-Aware Protection
Perceptual map. The key to soft restriction is an accurate map of perceptual sensitivity. We adopt just‑noticeable‑difference (JND) concept [65] because they are explicitly designed to predict the minimum perturbation that the human eye can detect. JND represents the minimum intensity of stimulus (perturbation in our context) required to produce a noticeable change in visual perception. The intent behind the JND estimation model is to determine this perceptual threshold. Given that the fundamental premise of JND— to quantify human sensitivity to subtle changes— aligns with our objective of perception-aware protection, we focus on analyzing the effectiveness of JNDs in our task, and which JNDs yield the good results.
To this end, as shown in Fig. 3, we compare five widely used JNDs: 1) luminance adaptation (LA), which considers pixel luminance [20], 2) contrast masking (CM), which measures scene complexity by luminance contrast [26, 63], 3) contrast sensitivity function (CSF), which utilizes frequency signals [62], 4) standard deviation (Stdev), which measures spatial structure with patch-wise statistics, and 5) entropy, which computes complexity with patch-wise entropy. More details are described in the Supplemental Materials.
In our observation, most JND models fit surprisingly well with soft restriction (Fig. 3, right). Consequently, we employ a JND-based perceptual maps, considering its simplicity and effectiveness. Note that among JNDs we used, LA and CM emerge as superior in the quality-protection trade-off. However, these results are the average scores across a dataset and we note that some artworks exhibit better trade-off with other JNDs. Especially, since artworks have diverse styles [12], the best combinations can differ; for example, in Fig. 3 (left), Stdev and entropy are best for top image while bottom image shows a preference for CM and CSF. Moreover, in practical scenarios, users who protect their artwork are solely concerned with their specific pieces, not the average scores. Thus, it is crucial to ensure that the protection method is effectively applied to every individual artwork.
@@
Instance-wise refinement. Although the above method is more effective in many scenarios than the single JND, applying a uniform averaged map across all artworks may still be suboptimal for some artworks since a single, dataset‑wide JND average cannot capture the stylistic diversity of real artworks. Moreover, the optimal perceptual map corresponding to an artwork varies depending on both artwork’s structure and the applied protective perturbations. Even for identical artwork, the detectability of perturbations is affected by changing these as human sensitivity varies with distortion type [11]. Considering these, we propose an instance-wise refinement (IWR), customizing the perceptual map $\mathcal{M}$ for each artwork. Instance‑wise refinement re‑weights the candidate JND maps per image, learning a convex combination that best matches that artwork’s own texture and color distribution while preserving PAP’s global imperceptibility objective. During optimization steps, the perceptual map is refined through a weighted sum: $\mathcal{M}(\omega)=\textstyle\sum_{k=1}^{K}\text{softmax}(\mathbf{\omega})^{k}*M^{k}$, where $\mathbf{\omega}$ is refinement parameters that adjust the contributions of each JND to $\mathcal{M}$. $\mathbf{\omega}$ is optimized using the objective function below:
[…]
@@
with $\mathcal{M}^{\prime}$ being the initial perceptual map before the refinement steps. In Eq. 7, the former term enforces the consistency of the refined perceptual map $\mathcal{M}$ with the initial map $\mathcal{M}^{\prime}$ by minimizing the discrepancy in protection loss between them. The latter term enhances the perception-aware protection for a given specific image. As shown in Fig. 3, instance-wise refinement (IWR) has a pronounced effect as compared to the naive averaging, better capturing the nuances of image-specific perceptual sensitivity.
Discussion. Restricting perturbation strength in a soft or gradual manner has been previously explored in the adversarial attack domain [40]. However, unlike adversarial attacks, which aim to fool recognition models, our task focuses on fooling generative models, a distinction that requires further analysis—something not addressed in existing work. In the image protection domain, similar soft restriction methods have been explored [53]. However, these methods did not aim to achieve more visually imperceptible protection. In addition, in our approach, we combine a perceptual map with instance-wise refinement, which, to the best of our knowledge, has not been explored in existing “soft” restriction methods. This combination is essential for robust style protection, ensuring that each artwork receives optimally distributed perturbations while maintaining imperceptibility for human viewers.
@@
Figure 4: Protection difficulty. We present the protection efficacy of PhotoGuard [50] across 20 artworks (x-axis). Notably, the level of style protection varies widely between artworks, thus the average score (red dashed line) is less meaningful to users who protect their artwork.
III-B Difficulty-Aware Protection
Motivation. While we adopt instance-wise refinement in PAP, it primarily focuses on “where” to apply protection, without explicitly specifying “how much” protection is needed. Therefore, it does not fully account for the image-specific nature. Empirically, we observe that different artworks exhibit varying levels of tolerance to perturbations before style leakage occurs. This motivates the need for a protection method that considers both where and how strongly to apply perturbations, while maintaining imperceptibility and effectiveness.
Figure 5: Perturbation amplification. Starting from protected images (left), the diffusion process (middle) amplifies imperceptible perturbations into visible distortions. These artifacts closely resemble those in the personalized outputs (right), suggesting that the diffusion trajectory exposes directions that personalization can exploit.
To this end, we additionally introduce difficulty-aware protection (DAP). Specifically, we observed that the difficulty of protection varies across images (or artworks); some images are protected properly with low protection strength, while others require higher strength to achieve the same protection level. Fig. 4 shows the protection efficacy of PhotoGuard [50] in 20 artworks. The score is calculated as $(\text{FID}_{\text{protected}}-\text{FID}_{\text{unprotected}})/\text{FID}_{\text{unprotected}}$, with a higher value indicating better protection efficacy. Surprisingly, the protection efficacy varies greatly across artworks. Moreover, some artworks show no difference between unprotected and protected cases. It implies that, for actual users, the average score might be a meaningless metric. So why does the level of protection differ between artworks? This question may tie into the learning dynamics of personalization methods, which are still unclear and require further investigation.
Nevertheless, we understand that the protection performance differs between artworks, so in practice, we can adjust the protection strength accordingly when we know the difficulty. The most straightforward way to predict the difficulty is to perform protection, personalize the protected/unprotected images with DreamBooth [48], and evaluate protection efficacy as done in Fig. 4. While effective, this process is extremely time-consuming, as it involves: (1) performing personalization on both clean (unprotected) and protected images, (2) comparing the results to evaluate the difficulty, and (3) reapplying protection based on the derived difficulty.
To make the above process more practical, we instead propose an approximated approach for difficulty-aware protection that utilizes the generative capacity of the same diffusion process but eliminating the need for a full personalization (i.e. DreamBooth fine-tuning). The core rationale stems from our preliminary empirical observations in Fig. 5: when applying forward and reverse diffusion processes to protected images, the injected perturbations tend to be amplified. Interestingly, these amplified perturbations closely resemble those exploited during DreamBooth-based personalization. In other words, if the diffusion process leads to significant perceptual changes, it implies that the corresponding regions in latent space are more vulnerable and can be easily exploited by personalization methods. Furthermore, prior work such as [67] has also shown that injecting perturbations into input images leads to amplification in the latent space ($z$-space), thereby indicating vulnerability. Similarly, our method reveals that even a partial diffusion process effectively exposes similar vulnerabilities. Building on this insight, we interpret such vulnerability as a proxy for protection difficulty, and accordingly adjust the protection strength based on the observed amplification during the diffusion process.